OpenStudy (darkigloo):

f(x)=x^4+x^2-2 On which interval is f increasing? A. (-1/√2, infinity ) B. (-1/√2 , 1/√2) C. (0, infinity) D. (-infinity, 0) E. (-infinity, -1/√2)

OpenStudy (anonymous):

take the derivative set it equal to zero, find intervals where is is positive and negative

OpenStudy (darkigloo):

i got x=0 and x= -1/√ 2 as critical points but i think thats wrong...

OpenStudy (anonymous):

you missed a zero

OpenStudy (anonymous):

oops it is a plus sign!~

OpenStudy (darkigloo):

so i was wrong?

OpenStudy (anonymous):

\[f'(x)=4x^3+2x=2x(x^2+1)\]

OpenStudy (anonymous):

\(x^2+1\) is never zero as \(x^2\geq 0\) always,so the only zero is at \(x=0\)

OpenStudy (darkigloo):

ohh ok thanks.